If you mix a 50% concentrate acic solution and a 20% concentrate acid solution and you end up with 40 gallons of a 42.5% concentrate acid solution, how much of the 50% and 20% solution did you use.
Answer by checkley71(8403)

.50X+.20(40-X)=.425*40
.5X+8-.2X=17
.3X=17-8
.3X=9
X=9/.3
X=30 GALLONS OF 50% ARE REQUIRED.
40-30=10 GALLONS OF 20% IS REQUIRED.
PROOF:
.5*30+.2*10=17
15+2=17
17=17